Solve the following initial value problems using method of undetermined coefficients: (ii) y''+y' = 5xsinx, y(0) = 0,y(0) = 1.


1
Expert's answer
2022-02-27T15:02:24-0500

"y''+y'=5xsinx\\
y(0)=0, y'(0)=1\\
y=y_1+y_2\\
1) y''+y'=0\\
\lambda^2+\lambda=0\\
\lambda(\lambda+1)=0\\
\lambda_1=0, \lambda_2=-1\\
y_1=C_1+C_2e^{-x}"

"2) y_2=(ax+b)cosx+(kx+p)sinx\\
y_2'=acosx-(ax+b)sinx+\\+ksinx+(kx+p)cosx=\\
=(kx+p+a)cosx+(-ax-b+k)sinx\\
y_2''=kcosx-(kx+p+a)sinx-\\-asinx+(-ax-b+k)cosx=\\
=(-ax-b+2k)cosx+(-kx-p-2a)sinx"


"(-ax-b+2k)cosx+(-kx-p-2a)sinx+\\
+(kx+p+a)cosx+(-ax-b+k)sinx=\\=5xsinx"


"xcosx: -a+k=0\\
cosx: -b+2k+p+a=0\\
xsinx:-k-a=5\\
sinx:-p-2a-b+k=0"


"a=-\frac{5}{2}, k=-\frac{5}{2}, b=\frac{5}{2}, p=5\\
y_2=(-\frac{5}{2}x+\frac{5}{2})cosx+(-\frac{5}{2}x+5)sinx"


"y=C_1+C_2e^{-x}+\\+(-\frac{5}{2}x+\frac{5}{2})cosx+(-\frac{5}{2}x+5)sinx\\
y'=-C_2e^{-x}-(-\frac{5}{2}x+\frac{5}{2})sinx-\\
-\frac{5}{2}cosx+(-\frac{5}{2}x+5)cosx-\frac{5}{2}sinx\\
y(0)=0\\
C_1+C_2+\frac{5}{2}=0\\
y'(0)=1\\
-C_2-\frac{5}{2}+5=1\\
C_2=\frac{3}{2}, C_1=-4"


"y=-4+\frac{3}{2}e^{-x}+\\+(-\frac{5}{2}x+\frac{5}{2})cosx+(-\frac{5}{2}x+5)sinx"